In ABC if b= 24 c= 20 sin(a/2)= 1/√8 find a
Please answer this question

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
Squaring on both sides 
Sn^2 (A/2) = 1/8 
CosA = 1-2Sin^2(A/2) 
CosA = 1-1/4 
CosA = 3/4 
A = Cos^-2(3/4)
Apply (b^2+c^2-a^2)/2bc = a-2/a 
Solve for a 
(976-a^2)/960  = (a-2)/a 
a^3 - 16a -1920 
Solve for a .
Hope this helps 
God Luck